# Can Graph Neural Networks Learn Language with Extremely Weak Text Supervision?

### v2

@@ Abstract
While great success has been achieved in building vision models with Contrastive Language-Image Pre-training (CLIP) over internet-scale image-text pairs, building transferable Graph Neural Networks (GNNs) with CLIP pipeline is challenging because of three fundamental issues: the scarcity of labeled data and text supervision, different levels of downstream tasks, and the conceptual gaps between domains. In this work, to address these issues, we leverage multi-modal prompt learning to effectively adapt pre-trained GNN to downstream tasks and data, given only a few semantically labeled samples, each with extremely weak text supervision. Our new paradigm embeds the graphs directly in the same space as the Large Language Models (LLMs) by learning both graph prompts and text prompts simultaneously. To accomplish this, we improve state-of-the-art graph prompt method, and then propose the first graph-language multi-modal prompt learning approach for exploiting the knowledge in pre-trained models. Notably, due to the insufficient supervision for fine-tuning, in our paradigm, the pre-trained GNN and the LLM are kept frozen, so the learnable parameters are much fewer than fine-tuning any pre-trained model. Through extensive experiments on real-world datasets, we demonstrate the superior performance of our paradigm in few-shot, multi-task-level, and cross-domain settings. Moreover, we build the first CLIP-style zero-shot classification prototype that can generalize GNNs to unseen classes with extremely weak text supervision.
Figure 1: CLIP backbone (top) and this work (bottom). If a research paper cites many papers from biology and computer science, we realize this paper will likely be about computational biology, even if we do not know what exactly computational biology is. CLIP builds image encoders that learn such language dependency by Contrastive Language-Image Pre-training in the same embedding space according to Internet-scale data. However, text supervision is often extremely weak for graphs. This work leverages Multi-modal Prompt Learning for Graph Neural Networks that can effectively teach GNNs language dependency given few training samples with weak text supervision.
@@ 1 Introduction
Graphs are constructed from real scenarios, but GNNs, optimized according to numerical labels, still do not understand what a label represents in the real world. To solve the issue of predetermined numerical categories, Contrastive Language-Image Pre-training (CLIP) [1] leverages natural language supervision by jointly training an image encoder and a text encoder in the same embedding space at scale. CLIP has demonstrated the ability to train high-quality, generalizable vision models [1, 2, 3, 4], which can adapt to diverse downstream tasks. Similar frameworks have been successfully extended to other modalities, including video [5, 6], 3D images [7, 8, 9], speech [10] and audio [11, 12], consistently demonstrating that alignment with text enhances the transferability of encoders. As for graphs, so far, such graph-text alignment has only been explored in the molecular domain [13, 14, 15, 16, 17] and on text-attributed graphs [18, 19, 20, 21, 22], where the paired graph-text data is relatively sufficient for pre-training.
@@
However, extending this paradigm to more general graph data poses significant challenges due to three facts. First, compared with language or vision data, graph data is very scarce and the text supervision is extremely weak [23, 24, 25] for pre-training models. Specifically, besides the number of samples being much smaller than images, many graph datasets are used for classification, where the label names consist of only a few tokens. Second, the task space of graph data could be on node-level [26, 27, 28, 29], edge-level [30, 31, 32], and graph-level [33, 34, 35, 36, 37, 38]. Third, in general, language tokens and visual objects retain the same conceptual meaning across different distributions, but the same graph structure may have distinct interpretations in different domains [39].
@@
Due to the above three reasons, jointly pre-training graph and text encoders is impractical for graph data with extremely weak text supervision. Fortunately, we can deal with the two modalities separately for pre-training: large language models have already been extensively pre-trained, and tremendous efforts have been devoted to pre-train GNNs through self-supervision [40, 41, 42, 43]. However, even with a pre-trained graph model, effectively adapting it to both the semantic embedding space for text alignment and diverse downstream tasks remains non-trivial. This raises a critical question:
@@
How to adapt pre-trained GNNs to the semantic embedding space given limited downstream data, i.e., few available samples and weak text supervision?
@@
This paper aims to answer this question based on the following observations: (1) Semantic text embedding spaces do not necessarily result from joint pre-training. In fact, the embedding spaces of encoder LLMs are inherently semantic and high-quality, as LLMs are trained on massive text data and demonstrate strong reasoning performance. (2) When the downstream data are limited, prompt learning [44, 45, 46, 47, 48] provides a better option than fine-tuning as much fewer parameters not only makes the optimization more efficient but also requires less resource than computing the gradient of a large model. Notably, some works have explored prompt learning for better alignment and obtained improvement in vision prediction [49, 50]. Inspired by these two observations, we propose a prompting-based paradigm with an LLM that aligns the GNN representations in the LLM’s semantic embedding space, while keeping the parameters of both GNN and LLM frozen.
@@
When attempting to adapt the representation from one modality to another, solely prompting a single modality could be sub-optimal, as it limits the adjustment to downstream tasks in the other modality [50]. To this end, we propose Multi-modal Prompt Learning for Graph Neural Networks (Morpher). Given a pre-trained GNN and few-shot semantically labeled graph data with weak text supervision, we assume zeroth-order access to a pre-trained LLM. Then, to leverage its high-quality semantic embedding space, Morpher connects and aligns the graph embeddings to it through prompting on both modalities with a cross-modal projector. Nonetheless, designing such a paradigm is more challenging than vision-language models. First, we lack jointly pre-trained encoders for the two modalities; instead, we only have two encoders whose embedding dimension is possibly different, pre-trained independently in each modality. Second, determining how to prompt the graph modality is non-trivial and remains a trending research topic. Third, the downstream data for GNN usually have much fewer labeled classes and labeled samples than V-L models, and the text supervision is extremely weak. Our contributions towards tackling these challenges are summarized as follows:
@@
  • Theoretically, we analyze that, in many cases, state-of-the-art graph prompt [51] is unable to learn good representations of the downstream data. We show that the optimization of the graph prompt is restricted by design. From the theoretical findings, we further improve state-of-the-art graph prompt according to the attention mechanism to prevent failure in optimization.
@@
  • To connect and adapt the pre-trained GNN with LLM effectively with extremely weak text supervision, we propose Morpher, a graph-text multi-modal prompt learning paradigm. To the best of our knowledge, this is the first approach to align the representations of GNN and LLM without fine-tuning any of their parameters.
@@
  • With extremely weak text supervision, we demonstrate the effectiveness of our improved graph prompt and Morpher on real-world datasets under few-shot, multi-task, and cross-domain settings. To show that GNN learns language dependency through Morpher, we present the first CLIP-style zero-shot generalization prototype where the GNN can correctly predict unseen classes.
@@ 2 Background
We use calligraphic letters (e.g., $\mathcal{A}$) for sets, and specifically $\mathcal{G}$ for graphs. We use bold capital letters for matrices (e.g., $\mathbf{A}$). For matrix indices, we use $\mathbf{A}(i,j)$ to denote the entry in the $i^{th}$ row and the $j^{th}$ column. Additionally, $\mathbf{A}(i,:)$ returns the $i^{th}$ row in $\mathbf{A}$.
@@
We use $\mathcal{G}=(\mathbf{A},\mathbf{X})$ to denote a graph with node set $\mathcal{V}$ and edge set $\mathcal{E}$, where $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ is the adjacency matrix and $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the node feature matrix. $\mathbf{A}(u,v)=1$ if there is an edge connecting $u$ and $v$; otherwise $\mathbf{A}(u,v)=0$. A Graph Neural Network $f_{\phi}^{g}(\cdot)$ with hidden dimension $d_{g}$ encodes $\mathcal{G}$ into the embedding space: $f_{\phi}^{g}(\mathcal{G})\in\mathbb{R}^{|\mathcal{V}|\times d_{g}}$, which could preserve both feature and structure information of $\mathcal{G}$. The extracted embeddings $f_{\phi}^{g}(\mathcal{G})$ can be used for various downstream tasks such as classification.
@@
Prompt learning adds learnable tokens to the downstream data and provides a powerful alternative to fine-tuning when the labeled downstream data is scarce. Prompt learning for encoders was first used in NLP. Let $f_{\phi}^{t}(\cdot)$ denote the LLM encoder with embedding dimension $d_{t}$. For a series of input tokens $\{x_{k}\}_{k=1}^{K}$, the LLM encoder embeds it as a matrix
$\mathbf{X}_{t}$ = $f_{\phi}^{t}(\{x_{k}\}_{k=1}^{K})\in\mathbb{R}^{K\times d_{t}}$, then aggregates the representation to a vector $aggre(\mathbf{X}_{t})\in\mathbb{R}^{1\times d_{t}}$ for downstream tasks. Prompt learning initializes a tunable matrix $\mathbf{P}_{\theta}^{t}\in\mathbb{R}^{n_{t}\times d_{t}}$, where $n_{t}$ denotes the number of text prompt tokens. Then, this tunable matrix is concatenated with the input tokens’ embeddings to form a single matrix $[\mathbf{P}_{\theta}^{t};\mathbf{X}_{t}]_{dim=0}\in\mathbb{R}^{(K+n_{t})\times d_{t}}$, and the aggregated vector for downstream tasks becomes $aggre([\mathbf{P}_{\theta}^{t};\mathbf{X}_{t}]_{dim=0})$. In practice, we can train the model to minimize the loss function for downstream tasks, with only the prompt parameters $\mathbf{P_{\theta}^{t}}$ being updated.
@@
Given a pre-trained GNN $f_{\phi}^{g}(\cdot)$ with embedding dimension $d_{g}$ and a pre-train LLM encoder $f_{\phi}^{t}(\cdot)$ with embedding dimension $d_{t}$. Without loss of generality, we assume the downstream task is graph-level classification, as we will show that the other types of GNN tasks can be reformulated as graph classification. For $L$-shot graph classification, we are given limited text-labeled pairs $\{(\mathcal{G}_{i},t_{c})\}_{i=1}^{L}$ for each class $c$. Each text label $t_{c}$ consists of only a few tokens. Assuming $\mathcal{T}$ is the set of all text labels $t_{c}$, we are provided a set of test graphs $\{\mathcal{G}_{j}\}_{j=1}^{L_{test}}$. Using the pre-trained GNN and LLM, we want to correctly predict the text label $t_{j}\in\mathcal{T}$ for each test graph $\mathcal{G}_{j}$.
3 Revisiting and Improving Prompt as Graphs
Unlike prompting text data (which appends learnable text tokens to the original text sequence) and prompting image data (which pads a learnable image area above the original image), prompting graph data presents a significant challenge due to the non-euclidean nature of graphs. The recent pioneering work [51] designs the graph prompt still as a graph, then inserts it into the original graph by computing the inner-connections within the prompt graph and the cross-connections between the prompt graph and the original graph. An advantage of prompting at the graph level is that the downstream tasks of GNN can be reformulated into graph-level tasks. For the node classification task, we can induce the $\gamma$-ego-graph of each node by extracting the subgraph within a pre-defined distance $\gamma$. Then, we treat the node label as the induced ego-graph label. Similarly, for the edge classification task, we can extract a subgraph for each edge by extending the node pair to their $\gamma$ distance neighborhood, and use the edge label as the induced graph label. By inducing subgraphs, we can reformulate node-level and edge-level downstream tasks to graph-level.
@@ 3 Revisiting and Improving Prompt as Graphs
To prompt a graph $\mathcal{G}$, each prompt token is a new node. Let $n_{g}$ denote the number of prompt tokens and $\mathcal{P}=\{p_{i}\}_{i=1}^{n_{g}}$ denote the set of prompt tokens. The graph prompt is formulated by a tunable matrix $\mathbf{P}_{\theta}^{g}\in\mathbb{R}^{n_{g}\times d}$, where $d$ is the node feature dimension of graph $\mathcal{G}$. In other words, each row vector $\mathbf{P}_{\theta}^{g}(i,:)$ is the feature of the prompt token $p_{i}$. Then, the mechanism to prompt a graph $\mathcal{G}=(\mathbf{A},\mathbf{X})$ with $n$ nodes and $d$ feature dimension is as follows [51].
  • Compute inner-connections to construct the prompt graph $\mathcal{G}_{p}=(\mathbf{A}_{p},\mathbf{X}_{p})$. For the feature matrix, we directly set $\mathbf{X}_{p}=\mathbf{P}_{\theta}^{g}$. For two prompt tokens $p_{i}$ and $p_{j}$, the prompt graph will have an edge between them if and only if the dot product of their features is larger than a threshold. In other words, $\mathbf{A}_{p}(i,j)=1\iff\sigma(\mathbf{P}_{\theta}^{g}(i,:)\mathbf{P}_{\theta}^{g}(j,:)^{\top})>\delta_{inner}$, where $\sigma(\cdot)$ is the sigmoid function.
  • Compute cross-connections to insert the prompt graph $\mathcal{G}_{p}$ into the original input graph $\mathcal{G}$. Similarly, for $x_{i}\in\mathcal{G}$ and $p_{j}\in\mathcal{G}_{p}$, there is an edge between them if and only if $\sigma(\mathbf{X}(i,:)\mathbf{P}_{\theta}^{g}(j,:)^{\top})>\delta_{cross}$.
@@
Since not all the GNN backbones can take edge weights [52], the cross-connections in a manipulated graph are discrete [^1] In official implementation of [51], adjacency matrices are discrete: either 0 or 1 for each entry., thresholded by $\delta_{cross}$. However, the input node features of most real-world datasets are sparse, resulting from the construction process [53, 54, 55]. As shown in Table 6, $||\mathbf{X}(i,:)||_{1}$ is typically 1. As the initialization of each token feature $\mathbf{P}_{\theta}^{g}(i,:)$ is close to $\vec{\boldsymbol{0}}$, for any node $i$ and token $p_{j}$, the dot products $\mathbf{X}(i,:)\mathbf{P}_{\theta}^{g}(j,:)^{\top}$ is close to 00 , and the sigmoid value is very close to $0.5$. Consequently, if we want the graph prompt to have cross-connections, we must set $\delta_{cross}<0.5$. Then, as the sigmoid values are close to 0.5, the cross-connections will be dense, i.e., almost every node in the original graph is connected with every node token in the prompt graph. For two different graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ in the same task, the prompt graph $\mathcal{G}_{p}$ is identical. Since the GNNs work by aggregating the node features, their embeddings $f_{\phi}^{g}(\mathcal{G}_{1})$ and $f_{\phi}^{g}(\mathcal{G}_{2})$ are approximately the same because the features in the prompt graph overwhelm the features in the original graphs due to the dense cross-connections. Then, according to a trivial lemma from optimization, even if $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ have different labels, the task head classifier cannot be trained to distinguish them [^2] In fact, when executing the official implementation of [51] on Cora, the training loss does not decrease. Similar problems have been observed by another work [56]..
Lemma 3.1.
For any classifier $c(\cdot)$, if the identical feature $\mathbf{x}$ has label distribution $p(\cdot)$, then the optimal classification for cross-entropy loss is $Pr(c(\mathbf{x})=y)=p(y)$. From this, if two graphs have similar embedding but different labels, GNN training may not converge. (Proof in Appendix A)
@@
The issue of the current graph prompt is rooted in the imbalance of original connections in the input graph and cross-connections between input and prompt, as shown in Figure 2 (left). We also visualize the standard NLP attention mechanism [57] in Figure 2 (middle). After we prepend a sequence of text prompt tokens $\{p_{i}^{t}\}$, the features of the text prompt tokens will be densely aggregated to the features of the original text tokens. In other words, the “cross-connection” between the text prompt sequence and the input sequence is dense. However, such a dense connection does not cause the prompt feature to overwhelm the input, because the features in the input sequence are also aggregated in a dense manner. Inspired by this, the number of cross-attention between input and prompt should approximate the number of input connections. Since the connection of a graph dataset is often sparse, we should also constrain the cross-connections between the prompt graph and the input graph to be sparse as well.
Nonetheless, “sparse” is a wide concept to implement: if the cross-connections are too dense, the prompt graph will dominate the input graph; but if the cross-attention is too sparse, the prompt graph will be limited to manipulating the input graph. We deem that a balance could be achieved by approximately equalizing the number of cross-connections with that of connections in the input graph, i.e., $n_{e}$. Therefore, we set the number of cross-connections to at most $n_{e}$ by connecting each node in the input graph with at most $\left\lfloor{\frac{n_{e}}{a}}\right\rfloor$ prompt tokens. Then, we can safely use a small $\delta_{cross}$ and cosine similarity $\frac{\mathbf{X}(i,:)\cdot\mathbf{P}_{\theta}^{g}(j,:)^{\top}}{\|\mathbf{X}(i,:)\|_{2}\|\mathbf{P}_{\theta}^{g}(j,:)\|_{2}}$ instead of $\sigma(\mathbf{X}(i,:)\mathbf{P}_{\theta}^{g}(j,:)^{\top})$ to calculate the cross-connections. We demonstrate that our improved graph prompt works better in the later experiments.
4 Multi-modal Prompt Learning for GNNs

### v3

@@ Abstract
While great success has been achieved in building vision models with Contrastive Language-Image Pre-training (CLIP) over internet-scale image-text pairs, building transferable Graph Neural Networks (GNNs) with CLIP pipeline is challenging because of the scarcity of labeled data and text supervision, different levels of downstream tasks, and the conceptual gaps between domains. In this work, to address these issues, we propose a multi-modal prompt learning paradigm to effectively adapt pre-trained GNN to downstream tasks and data, given only a few semantically labeled samples, each with extremely weak text supervision. Our new paradigm embeds the graphs directly in the same space as the Large Language Models (LLMs) by learning both graph prompts and text prompts simultaneously.dai We demonstrate the superior performance of our paradigm in few-shot, multi-task-level, and cross-domain settings. Moreover, we build the first CLIP-style zero-shot classification prototype that can generalize GNNs to unseen classes with extremely weak text supervision. The code is available at https://github.com/Violet24K/Morpher.
Can Graph Neural Networks Learn Language
with Extremely Weak Text Supervision?
Zihao Li1, Lecheng Zheng1, Bowen Jin1, Dongqi Fu2, Baoyu Jing1, Yikun Ban1, Jingrui He1, Jiawei Han1 1University of Illinois Urbana-Champaign, 2Meta AI Correspondence: zihaoli5@illinois.edu
@@ 1 Introduction
Graphs are constructed from real scenarios, but GNNs, optimized according to numerical labels, still do not understand what a label represents in the real world. To solve the issue of predetermined numerical categories, CLIP (Radford et al., 2021) leverages natural language supervision by jointly training an image encoder and a text encoder in the same embedding space at scale. CLIP has demonstrated the ability to train high-quality, generalizable vision models (Radford et al., 2021; Jia et al., 2021; Li et al., 2022), which can adapt to diverse downstream tasks. Similar frameworks have been successfully extended to video (Xu et al., 2021), 3D images (Hess et al., 2024), speech (Shih et al., 2022) and audio (Guzhov et al., 2022), consistently demonstrating that alignment with text enhances the transferability of encoders. As for graphs, so far, such graph-text alignment has only been explored in the molecular domain (Luo et al., 2023; Liu et al., 2023e) and on text-attributed graphs (Wen and Fang, 2023; Li et al., 2023a; Jin et al., 2023b; Yan et al., 2023), where the paired graph-text data is relatively sufficient for joint pre-training.
@@
However, extending this paradigm to more general graph data poses significant challenges due to three facts. First, compared with language or vision data, graph data is very scarce and the text supervision is extremely weak (Liu et al., 2023a; Chen et al., 2023b; Manchanda et al., 2023). Besides the number of samples being much smaller than images, many graph datasets are used for classification, where the label names consist of only a few tokens. Second, the task space of graph data could be on node-level, edge-level, and graph-level. Third, in general, language tokens and visual objects retain the same conceptual meaning across different distributions, but the same graph structure may have distinct interpretations in different domains.
@@
Jointly pre-training graph and text encoders is impractical for graph data with extremely weak text supervision. Fortunately, we can deal with the two modalities separately for pre-training: large language models have already been extensively pre-trained, and tremendous efforts have been devoted to pre-train GNNs through self-supervision (Hu et al., 2020a; Liu et al., 2023c; Zheng et al., 2024c, d). However, even with a pre-trained graph model, effectively adapting it to both the semantic embedding space for text alignment and diverse downstream tasks remains non-trivial. This raises a critical question:
@@
How to adapt pre-trained GNNs to the semantic embedding space given limited downstream data, i.e., few samples and weak text supervision?
@@
This paper aims to answer this question based on the following observations: (1) Semantic text embedding spaces do not necessarily result from joint pre-training. In fact, the embedding spaces of encoder LLMs are inherently semantic and high-quality, as LLMs are trained on massive text data and demonstrate strong performances. (2) When the downstream data are limited, prompt learning (Li and Liang, 2021; Houlsby et al., 2019; Zhang et al., 2022a; Lester et al., 2021) provides a better option than fine-tuning as much fewer parameters not only makes the optimization more efficient but also requires less resource than fine-tuning a large model. Notably, some works have explored prompt learning for better alignment and obtained improvement in vision prediction (Zhou et al., 2022b; Khattak et al., 2023). Inspired by these observations, we propose a prompting-based paradigm with an LLM that aligns the GNN representations in the semantic embedding space, while keeping the parameters of both GNN and LLM frozen.
@@
When adapting the representation from one modality to another, solely prompting a single modality could be sub-optimal, as it limits the adjustment to downstream tasks in the other modality (Khattak et al., 2023). To this end, we propose Multi-modal Prompt Learning for Graph Neural Networks (Morpher). Given a pre-trained GNN and few-shot semantically labeled graph data with weak text supervision, we assume zeroth-order access to a pre-trained LLM. Then, to leverage its high-quality semantic embedding space, Morpher connects and aligns the graph embeddings to it through prompting on both modalities with a cross-modal projector. Nonetheless, designing such a paradigm is more challenging than vision-language models. First, we lack jointly pre-trained encoders for the two modalities; instead, we only have two encoders pre-trained independently in each modality. Second, determining how to prompt the graph modality is non-trivial and remains a trending research topic. Third, the downstream data for GNN usually have much fewer labeled classes and labeled samples than Vision-Language models, and the text supervision is extremely weak. Our contributions towards tackling these challenges are:
@@
  • We analyze that, state-of-the-art graph prompt (Sun et al., 2023a) is often unable to learn good representations of the downstream data. We further improve it to prevent unstable optimization.
@@
  • To connect and adapt the pre-trained GNN with LLM effectively with extremely weak text supervision, we propose Morpher, the first graph-text multi-modal prompt learning paradigm to align the representations of GNN and LLM without fine-tuning any of their parameters.
@@
  • With extremely weak text supervision, we demonstrate our improved graph prompt and Morpher under few-shot, multi-task, and cross-domain settings. To show that GNN learns language dependency through Morpher, we present the first CLIP-style zero-shot generalization prototype where the GNN can predict unseen classes.
@@ 2 Background
We use calligraphic letters (e.g., $\mathcal{A}$) for sets, and specifically $\mathcal{G}$ for graphs. We use bold capital letters for matrices (e.g., $\mathbf{A}$). For matrix indices, we use $\mathbf{A}(i,j)$ to denote the entry in the $i^{th}$ row and the $j^{th}$ column. $\mathbf{A}(i,:)$ is the $i^{th}$ row in $\mathbf{A}$.
@@
We use $\mathcal{G}=(\mathbf{A},\mathbf{X})$ to denote a graph with node set $\mathcal{V}$ and edge set $\mathcal{E}$, where $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ is the adjacency matrix and $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the node feature matrix. $\mathbf{A}(u,v)=1$ if there is an edge connecting $u$ and $v$; otherwise $\mathbf{A}(u,v)=0$. A graph neural network $f_{\phi}^{g}(\cdot)$ with hidden dimension $d_{g}$ encodes $\mathcal{G}$ into the embedding space: $f_{\phi}^{g}(\mathcal{G})\in\mathbb{R}^{|\mathcal{V}|\times d_{g}}$, which could preserve both feature and structure information of $\mathcal{G}$.
@@
Let $f_{\phi}^{t}(\cdot)$ denote the LLM encoder with embedding dimension $d_{t}$. For a series of input tokens $\{x_{k}\}_{k=1}^{K}$, the LLM encoder embeds it as a matrix $\mathbf{X}_{t}$ = $f_{\phi}^{t}(\{x_{k}\}_{k=1}^{K})\in\mathbb{R}^{K\times d_{t}}$. Prompt learning initializes a tunable matrix $\mathbf{P}_{\theta}^{t}\in\mathbb{R}^{n_{t}\times d_{t}}$, where $n_{t}$ denotes the number of text prompt tokens. Then, this tunable matrix is concatenated with the input tokens’ embeddings to form $[\mathbf{P}_{\theta}^{t};\mathbf{X}_{t}]_{dim=0}\in\mathbb{R}^{(K+n_{t})\times d_{t}}$.
@@
Given a pre-trained GNN $f_{\phi}^{g}(\cdot)$ with embedding dimension $d_{g}$ and a pre-trained LLM encoder $f_{\phi}^{t}(\cdot)$ with embedding dimension $d_{t}$. Without loss of generality, we assume the downstream task is graph-level classification, as node-level or edge-level GNN tasks can be reformulated as graph-level by inducing ego-graphs within neighbor distance $\gamma$. For $L$-shot graph classification, we are given limited text-labeled pairs $\{(\mathcal{G}_{i},t_{c})\}_{i=1}^{L}$ for each class $c$. Each text label $t_{c}$ consists of only a few tokens. Assuming $\mathcal{T}$ is the set of all text labels $t_{c}$, we are provided a set of test graphs $\{\mathcal{G}_{j}\}_{j=1}^{L_{test}}$ and want to predict the text label $t_{j}\in\mathcal{T}$ for each test graph $\mathcal{G}_{j}$.
3 Improving Single-modal Graph Prompt
Unlike prompting text data, prompting graph data presents a significant challenge due to the non-euclidean nature of graphs. The pioneering work (Sun et al., 2023a) designs the graph prompt still as a graph and then inserts it into the original graph.
@@ 3 Improving Single-modal Graph Prompt
To prompt a graph $\mathcal{G}$, each prompt token is a new node. Let $n_{g}$ denote the number of prompt tokens and $\mathcal{P}=\{p_{i}\}_{i=1}^{n_{g}}$ denote the set of prompt tokens. The graph prompt is formulated by a tunable matrix $\mathbf{P}_{\theta}^{g}\in\mathbb{R}^{n_{g}\times d}$, where $d$ is the node feature dimension. Each row vector $\mathbf{P}_{\theta}^{g}(i,:)$ is the feature of the prompt token $p_{i}$. Then, the mechanism to prompt a graph $\mathcal{G}=(\mathbf{A},\mathbf{X})$ with $n$ nodes and $d$ feature dimension is Sun et al. (2023a)
  • Compute inner-connections to construct the prompt graph $\mathcal{G}_{p}=(\mathbf{A}_{p},\mathbf{X}_{p})=(\mathbf{A}_{p},\mathbf{P}_{\theta}^{g})$. For two prompt tokens $p_{i}$ and $p_{j}$, $\mathbf{A}_{p}(i,j)=1\iff\sigma(\mathbf{P}_{\theta}^{g}(i,:)\mathbf{P}_{\theta}^{g}(j,:)^{\top})>\delta_{inner}$, where $\sigma(\cdot)$ is the sigmoid function.
  • Compute cross-connections to insert the prompt graph $\mathcal{G}_{p}$ into $\mathcal{G}$. Similarly, for $x_{i}\in\mathcal{G}$ and $p_{j}\in\mathcal{G}_{p}$, there is an edge between them if and only if $\sigma(\mathbf{X}(i,:)\mathbf{P}_{\theta}^{g}(j,:)^{\top})>\delta_{cross}$.
@@
The input node features of most real-world datasets are sparse, resulting from the construction process (Yang et al., 2016; Morris et al., 2020; Dwivedi et al., 2023). As shown in Appendix Table 4, $||\mathbf{X}(i,:)||_{1}$ is typically 1. As the initialization of each token feature tensor $\mathbf{P}_{\theta}^{g}(i,:)$ is close to $\vec{\boldsymbol{0}}$ to stabilize gradients, for any node $i$ and token $p_{j}$, the dot products $\mathbf{X}(i,:)\mathbf{P}_{\theta}^{g}(j,:)^{\top}$ is close to 00, and the sigmoid value is very close to $0.5$. Consequently, if we want the graph prompt to have cross-connections, we must set $\delta_{cross}<0.5$. Then, as the sigmoid values are close to 0.5, the cross-connections will be dense, i.e., almost every node in the original graph is connected with every node token in the prompt graph. For two different graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ in the same task, the prompt graph $\mathcal{G}_{p}$ is identical. Since the GNNs aggregate the node features, their embeddings $f_{\phi}^{g}(\mathcal{G}_{1})$ and $f_{\phi}^{g}(\mathcal{G}_{2})$ are approximately the same because the features in the prompt graph overwhelm the features in the original graphs due to the dense cross-connections. Then, even if $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ have different labels, the task head classifier cannot be trained to distinguish them  [^1] In fact, similar training instability problems have been observed by another work (Zhao et al., 2024)..
In Appendix C.1, we show that initializing graph prompt token feature tensor with higher variance cannot effectively address this problem.
@@
The issue of the current graph prompt lies in the significant imbalance between original connections within the input graph and the input-prompt cross-connections, as illustrated in Figure 1 (left). We also visualize the standard NLP attention mechanism (Vaswani et al., 2017) in Figure 1 (middle). When a sequence of text prompt tokens ${p_{i}^{t}}$ is prepended to the input sequence, the features of the prompt tokens will be aggregated with those of the input tokens through a dense “cross-connection”, i.e., attention. Simultaneously, the features within the input sequence are also densely aggregated via attention, maintaining a balance with the prompt-input aggregation to prevent overwhelming. Inspired by this, we deem that a balance could be achieved by approximately equalizing the number of cross-connections with that of input graph connections, i.e., $n_{e}$. Since the connection of a graph dataset is often sparse, we constrain the cross-connections to be sparse as well. Therefore, we set the number of cross-connections to at most $n_{e}$ by connecting each node in the input graph with at most $\left\lfloor{\frac{n_{e}}{a}}\right\rfloor$ prompt tokens. Then, we can safely use a small $\delta_{cross}$ and cosine similarity $\frac{\mathbf{X}(i,:)\cdot\mathbf{P}_{\theta}^{g}(j,:)^{\top}}{\|\mathbf{X}(i,:)\|_{2}\|\mathbf{P}_{\theta}^{g}(j,:)\|_{2}}$ instead of $\sigma(\mathbf{X}(i,:)\mathbf{P}_{\theta}^{g}(j,:)^{\top})$ to calculate the cross-connections.
4 GNN Multi-modal Prompt Learning